OpenStudy (anonymous):

A wire runs from the top of a flagpole to a point on the ground that is 9 yards from the base of the pole. The wire is 15 yards long. How many yards tall is the flagpole? (Please type ONLY the numerical answer, no extra spaces or words. Do NOT include any units.)

OpenStudy (lgbasallote):

 use pythagorean theorem

OpenStudy (lgbasallote):

\[b = \sqrt{c^2 - a^2}\]

OpenStudy (lgbasallote):

\[b = \sqrt{15^2 - 9^2}\]

OpenStudy (lgbasallote):

\[b = \sqrt{225 - 81}\]

OpenStudy (lgbasallote):

\[b = \sqrt{144}\] \[b = 12\]

OpenStudy (anonymous):

Thank you.